1/2 = 8m
I have no idea what to do.

Answer by Alan3354(69443)

1/2 = 8m
I have no idea what to do.
=============
Step 1, say what you want to do.
If you want to solve for m:
1/2 = 8m
Divide both sides by 8
1/16 = m
m = 1/16